OpenStudy (anonymous):

The solution of an absolute-value equation is given. What is the equation? x= 2 + or - 3

OpenStudy (anonymous):

x = 2 + 3 = 5 x = 5 x-5 = 0 x= 2-3 x = -1 x + 1 = 0 I think this is what you are asking?

OpenStudy (anonymous):

start with each of your solutions... 2+ 3 and 2-3

OpenStudy (anonymous):

Hmm I think you're only suppose to get one equation overall.

OpenStudy (anonymous):

With absolute value brackets!

OpenStudy (anonymous):



OpenStudy (anonymous):

Alrighty thanks!